# Empirical investigation of multi-source cross-validation in clinical machine learning

### v1

@@ 5 Discussion
The experiments comparing stratified K-fold CV and LSO CV revealed the expected result that models trained in one context, may exhibit poor performance when used on a new data source. Comprehending the reasons behind this phenomenon is particularly crucial when these models are intended for use in healthcare diagnostics and monitoring to avoid challenges in deployment. They can involve investigating variations in patient demographics and data collection or diagnostic practices across different data sources.
Additionally, differences in data source sizes and label prevalence put more emphasis on certain labels, while differences in labeling practices can influence the model’s ability to generalize to unseen data. Of the data sources, the CPSC and CPSC-Extra dataset exhibit a high degree of variety in label proportions (Table 3) and showed low test performance in both experimental settings. Moreover, since the sources used for model training are not directly from the same distribution as the test set, the generalization power of the models may be reduced, leading to overoptimistic prediction performance in the case of K-fold CV. While it may not always be needed to train highly generalizable models [39], understanding the disparities between the training context and a test site is essential for reliable model performance evaluation and deployment considerations.
@@
Such multi-source experiments are only possible if multiple data sources are available and they may give a reliable picture on the performance range where the model on unseen data would lead. The main benefit in such multi-source experiment is that they provide good insights about the uncertainty levels of the prediction performance. In the development of predictive and decision-support models to healthcare, it’s crucial to understand how properties of the data used to train a model can can easily lead to overoptimistic expectations of the model’s actual performance and thus, potentially leading to a failure in deployment. However, to further understand this effect that’s a result of different training and test contexts, domain adaptation and transfer learning could be considered in the future studies.
@@
One of the main limitations in our study is that most of the data sources – the CPSC and CPSC-Extra dataset, the Chapman-Shaoxing and Ningbo dataset and the SPH data – are gathered from one country, namely China, covering over 70 % of the ECGs recordings. Despite the significance of these sources, the geographical centralization may introduce bias and limit the generalizability of models to hospitals from other countries. Within medical data, specific concerns have raised regarding the use of non-diverse and non-representative data, particularly when it inadequately represents the diversity of population, which can result in biased algorithms and, consequently, less accurate performance [40].
@@
Additionally, we only examined the generalization ability of ResNet [41] models and we did not evaluate different augmentations of ECGs, which are clear limitations for our study. Data augmentation in ECG classification has demonstrated potential to tackle with label imbalance, e.g. in forms of an amplitude alternation [42] and generative adversarial networks [43]. Also, albeit the success of the convolutional neural networks like ResNets for time series classification is evident, other network architectures such as a long short-term memory networks have shown their potential [44, 45]. These are some additional factors to consider in the future studies to further develop ECG classifiers.
While cross-validation methods are valuable tools to evaluate developed models, the context to which they are employed should be carefully considered. Researchers should stay vigilant in assessing whether the chosen approach genuinely addresses the specific research question at hand. It is crucial to avoid drawing misleading conclusions about a model’s actual abilities in order to advance the deployment of machine learning models tailored for modern healthcare applications.
@@ 6 Conclusion
In this paper we evaluated the reliability of the estimates of two CV methods, the standard K-fold CV and the LSO CV. To examine the generalization of ECG based classifiers in multi-source context, we created a large multi-source resource by combining and harmonizing two openly available but separate datasets, the PhysioNet/CinC challenge 2021 training data and the Shandong Provincial Hospital database. We performed two sets of experiments in single and multi-source settings, a ”source” corresponding to different ECG datasets. We showed that the K-fold CV provides overoptimistic performance claims of the model’s generalizability power in both settings, In single source experiment, the MEs of the 5-fold CV are from between 0.015–0.096 (macro) and 0.047–0.174 (micro). In multi-source experiment, the ME of the LSO CV, is close to zero, but the ME of the 4-fold CV is approximately 0.04 AUC. For future studies, the enhancement of deep learning model’s generalization can be examined through the utilization of data augmentation techniques and other model architectures. Additionally, to further study the effect when a model trained in one context performs poorly in another context, data adaptation and transfer learning could be considered.
7 Code availability
The source code that was used to perform the experiments can be found at https://github.com/tuijalei/ecg_analysis. The repository is a refactored version of the repository by Zhao et al. [8]. It contains functionalities to i) save demographic information, diagnostic labels and ECG paths to a CSV file which is used during model training and evaluation, ii) store parameters and other training and validation settings for each run as YAML files, iii) map the AHA codes to the SNOMED CT codes and vice versa, iv) preprocess ECGs using different transformations v) and train deep learning models and make predictions using trained models.
8 Acknowledgments
The authors wish to acknowledge CSC – IT Center for Science, Finland, for generous computational resources. This work has been supported by Research Council of Finland (grants 352893, 358868).

### v2

@@ 5 Discussion
The experiments comparing stratified K-fold CV and LSO CV revealed that models trained in one context may exhibit poor performance on a new data source. While this issue is well-known in research [10, 21, 17, 18], it has not been extensively analyzed within medical data. A major contributing factor behind the issue has been the lack of sufficiently diverse datasets [15, 16]. However, even with the availability of data, its characteristics play a significant role in model performance and generalizability. Within medical data, there is considerable variability in patient demographics, clinical settings, and data collection methods, all of which can influence how well a model performs on new context and highlights the importance of evaluating the robustness of predictive models — whether based on deep learning or classical techniques — across diverse datasets. Thus, it is essential not only to apply CV methods adequately but also to account for the data’s characteristics and variability, as these factors influence the effectiveness of the methods. This is especially important to avoid over-optimistic performance claims (e.g., [19]). Although training highly generalizable models may not always be necessary in every scenario [43], understanding the differences between the training context and a test site remains vital for reliable model performance evaluation and deployment in healthcare applications.
Building on these challenges, issues such as varying labeling standards, data bias, and ambiguous ground truth are evident [6, 7, 16, 9], and they remain relevant to our study. When harmonizing data from multiple sources, the inconsistencies can significantly impact model performance, and instead of learning physiological patterns in data, models may inadvertently learn the noise specific to each source. For example, we wanted to include the SR label for the SPH data as such label was included to the CinC data and otherwise, the SPH data would only contain abnormal ECGs. Although we verified the label mapping between SNOMED CT and AHA with the cardiologists, differences in labeling practices between the sources are expected: As PhysioNet also noted, they did not correct for variations in label practices - whether machine-generated or cardiologist-provided - within their data [12]. Our experiments using single-source data demonstrated that models trained on either of the CinC data sources performed well with the SPH data, yet displayed increased uncertainty when applied to other CinC sources, which may suggest potential biases and differences between the data sources. Additionally, variation in data source sizes (Table 1) and label prevalence (Table 3 and Figure 1) put more emphasis on certain labels, which may skew model generalization and is often observed in medical data. Of the data sources, the CPSC and CPSC-Extra source exhibit a high degree of variety in label proportions and showed low test performance especially in single source experiments. Moreover, labeling is not the only factor distinguishing the data sources; age and sex distributions (Table 2), as well as recording conditions and instrumentation, also contribute to differences. When models are trained on data from one source, they may inadvertently learn patterns or noise specific to that source, rather than capturing generalizable physiological patterns from ECGs.
@@
Such multi-source experiments are only possible if multiple data sources are available and they may give a reliable picture on the performance range where the model on unseen data would lead. The main benefit in such multi-source experiment is that they provide good insights about the uncertainty levels of the prediction performance particularly when the objective is to develop models for application to new sources, such as hospitals where training data may not be available. In such context, the LSO CV can be particularly effective in providing a clearer picture of the model’s actual performance on unseen data. In the development of predictive and decision-support models to healthcare, it’s crucial to understand how properties of the data used to train a model can easily lead to overoptimistic expectations of the model’s actual performance and thus, potentially leading to a failure in deployment. However, to further understand this effect that’s a result of different training and test contexts, domain adaptation and transfer learning could be considered in the future studies.
@@
One of the main limitations in our study is that most of the data sources – the CPSC and CPSC-Extra source, the Chapman-Shaoxing and Ningbo source and the SPH data – are gathered from one country, namely China, covering over 70 % of the ECGs recordings. Despite the significance of these sources, the geographical centralization may introduce bias and limit the generalizability of models to hospitals from other countries. Within medical data, specific concerns have raised regarding the use of non-diverse and non-representative data, particularly when it inadequately represents the diversity of population, which can result in biased algorithms and, consequently, less accurate performance [44]. Although we were able to identify and address duplicate recordings in the SPH source, we could not do the same for the CinC data. Consequently, duplicate entries were retained in the final dataset, as these duplicates constitute only a very small fraction of the total data and are unlikely to have a significant impact on the results. However, this is a challenge that must be acknowledged when working with medical datasets, as it can affect claims regarding model performance [19].
@@
Additionally, we did not evaluate different augmentations of ECGs nor tuned hyper-parameters for the neural networks, which are clear limitations for our study. Data augmentation in ECG classification has demonstrated potential to tackle with label imbalance, e.g. in forms of an amplitude alternation [45] and generative adversarial networks [46]. Also, albeit the success of the convolutional neural networks like ResNets for time series classification is evident, other network architectures such as a long short-term memory networks have shown their potential [47, 48]. These are some additional factors to consider in the future studies to further develop ECG classifiers and study their capabilities in CVD classification.
While cross-validation methods are valuable tools to evaluate developed models, the context to which they are employed should be carefully considered. Researchers should stay vigilant in assessing whether the chosen approach genuinely addresses the specific research question at hand. The medical data is inherently complex, with variability stemming from differences in patient populations, data collection practices, and clinical environments, which can add challenges in model evaluation. As we demonstrated, the estimated performance of a model can vary significantly between different CV methods, albeit each method addresses different aspects of the model’s generalization capabilities. For example, the K-fold CV is suitable when multiple hospitals collaborate to develop a model for use within those same hospitals. However, when the goal is to develop a model for a hospital from which training data is no available, the LSO CV may provide more reliable performance estimates. A limitation of LSO CV is that if the number of sources is small, the performance estimates obtained can be pessimistic. For example, if there is data from only two or three hospitals, removal of one of these could significantly reduce the variety of training data and hence accuracy of trained models. Yet, as also highlighted by previous studies underscoring the importance of external validation on diverse yet relevant patient populations [20], it can be argued that it is better to err on the side of caution than provide overoptimistic evaluations when evaluating clinical prediction models.
Regardless of the CV method, it is advisable to assess the presence of batch effects in the data as a standard practice, as demonstrated by our approach using source-based classification. It is crucial to avoid drawing misleading conclusions about a model’s actual abilities in order to advance the deployment of machine learning models tailored for modern healthcare applications. In addition to performance evaluation, the LSO CV can also be used for model selection in order to choose hyperparameters, features or classifier algorithms likely to generalize well to new sources. If both model selection and final evaluation are to be performed simultaneously, nested [49] LSO CV could be implemented, though at considerable computational cost.
@@ 6 Conclusion
In this paper we evaluated the reliability of the estimates of two CV methods, the standard K-fold CV and the LSO CV. To examine the generalization of ECG-based classifiers in multi-source context, we created a large multi-source resource by combining and harmonizing two openly available but separate datasets, the PhysioNet/CinC challenge 2021 training data and the Shandong Provincial Hospital database. We conducted a source-based classification to analyze potential biases between the data sources. For CV, we performed two sets of experiments in single and multi-source settings, a ”source” corresponding to different ECG datasets. We showed that the K-fold CV provides overoptimistic performance claims of the model’s generalizability power in both settings. For future studies, the enhancement of deep learning model’s generalization can be examined through the utilization of data augmentation techniques and other model architectures. Additionally, to further study the effect when a model trained in one context performs poorly in another context, data adaptation and transfer learning could be considered.
Code availability
The source code that was used to perform the experiments can be found at https://github.com/UTU-Health-Research/dl-ecg-classifier. It contains functionalities to 1) save demographic information, diagnostic labels and ECG paths to a CSV file which is used during model training and evaluation, 2) store parameters and other training and validation settings for each run as YAML files, 3) map the AHA codes to the SNOMED CT codes and vice versa, 4) preprocess ECGs using different transformations, 5) and train and run models.
CRediT authorship contribution statement
Tuija Leinonen: Software, Data Curation, Formal analysis, Investigation, Writing - Original Draft, Visualization. David Wong: Methodology, Writing - Review & Editing. Antti Vasankari: Software & Writing - Original Draft. Ali Wahab: Resources, Data Curation. Ramesh Nadarajah: Resources, Data Curation. Matti Kaisti: Supervision, Conceptualization, Methodology, Formal analysis, Writing - Review & Editing, Funding acquisition. Antti Airola: Supervision, Conceptualization, Methodology, Formal analysis, Writing - Review & Editing, Funding acquisition.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
The authors wish to acknowledge CSC – IT Center for Science, Finland, for generous computational resources. This work has been supported by Research Council of Finland (grants 352893, 358868, 345805, 340140).